OpenStudy (anonymous):

The Caddo County Coop uses the function S(t) = -5,250t + 57,750 to determine the salvage value S(t), in dollars, of a fertilizer blender t years after its purchase. How long will it take the blender to depreciate completely?

OpenStudy (anonymous):

answer options 9 11 12 13

OpenStudy (anonymous):

to depreciate completely means 0 = -5,250t + 57,750 so the answer is 11

OpenStudy (anonymous):

to depreciate means 0>-5,250t + 57,750

OpenStudy (anonymous):

oh okay, would this be done like that oo? A forest owned by the Jumbo Lumber Corporation contains 4,288 trees. If Jumbo Lumber cuts down 75 trees every day, which function can be used to find the number of trees, n, left in the forest after t days?

OpenStudy (anonymous):

n(t) = (4,288-75)t n(t) = 4,288 - 75t n(t) = 4,288 + 75t n(t) = 4,288t - 75 ^ answer options

OpenStudy (anonymous):

of course B

OpenStudy (anonymous):

4,288-75t

OpenStudy (anonymous):

thanks :)